# Statistical Inference in Reinforcement Learning: A Selective Survey

### v1

@@ Abstract
Reinforcement learning (RL) is concerned with how intelligence agents take actions in a given environment to maximize the cumulative reward they receive. In healthcare, applying RL algorithms could assist patients in improving their health status. In ride-sharing platforms, applying RL algorithms could increase drivers’ income and customer satisfaction. Over the past decade, RL has been arguably one of the most vibrant research frontiers in machine learning. Nevertheless, statistics as a field, as opposed to computer science, has only recently begun to engage with RL both in depth and in breadth. This paper present a selective review of statistical inferential tools for RL, covering both hypothesis testing and confidence interval construction. Our goal is to highlight the value of statistical inference in RL for both the statistics and machine learning communities, and to promote the broader application of classical statistical inference tools in this vibrant area of research.
keywords:
Reinforcement learning; Statistical inference; Hypothesis testing; Confidence interval.
@@ 1 Introduction
In the statistics literature, RL is closely related to a vast body of work on estimating dynamic treatment regimes (DTRs), which aims to identify personalized treatment strategies for individual patients to maximize their health outcomes; see Chakraborty and Moodie (2013); Kosorok and Moodie (2015); Kosorok and Laber (2019); Tsiatis et al. (2019); Li et al. (2023) for reviews. However, most methods employ the so-called non-Markov decision process (Kallus and Uehara, 2020) – where rewards and transitions depend on the entire data history – to model the observed data. They often prove ineffective when applied to Markov decision processes (Kallus and Uehara, 2022) – which are more commonly studied in machine learning to formulate numerous sequential decision making problems. More recent work has shifted its focus to studying Markov decision processes for both policy optimization (Ertefaie and Strawderman, 2018; Luckett et al., 2020; Liao et al., 2022; Yang et al., 2022; Chen et al., 2024; Li et al., 2024; Shi et al., 2024; Zhou et al., 2024) and policy evaluation (Liao et al., 2021; Hu and Wager, 2023; Ramprasad et al., 2023; Wang et al., 2023; Shi et al., 2024).
@@ 2 Background: Data, Policies, and Models in Reinforcement Learning
  • At each time step $t$, an agent observes certain features of the environment, denoted as ${\color[rgb]{1,.5,0}\bm{O}_{t}}$;
  • Based on ${\color[rgb]{1,.5,0}\bm{O}_{t}}$, the agent takes an action, denoted as ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$;
  • The environment provides a corresponding reward, denoted as ${\color[rgb]{1,0,0}\bm{R}_{t}}$, and transitions to a new state at the next time step.
This process repeats iteratively, as illustrated in Figure 1(a). Thus, the observed data can be summarized into a sequence of “observation-action-reward” triplets $({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}},{\color[rgb]{1,0,0}\bm{R}_{t}})_{t\geq 0}$. It is worth noting that the observation ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ at each time step is not equivalent to the environment’s state ${\color[rgb]{0,0,1}\bm{S}_{t}}$. Indeed, the state can be viewed as a special observation with the Markov property, and we will elaborate on the difference between the two later.
@@
    History-dependent policy: This is the most general form of policy. At each time $t$, we define $\bm{H}_{t}$ as the set containing the current observation ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ and all prior historical information $({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\bm{R}_{i}})_{i<t}$. Under this policy, the selection of ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$ depends on the entire history $\bm{H}_{t}$.
@@
    Markov policy: The second type is a subset of the first. Under this policy, the conditional distribution of ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$ depends only on the current observation ${\color[rgb]{1,.5,0}\bm{O}_{t}}$, rather than the entire history $\bm{H}_{t}$. In other words, for any $t$ and $a$, we have $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a|\bm{H}_{t})=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\bm{O}_{t}})$.
@@
    Stationary policy: The third type is a subset of the second. Under this policy, the conditional distribution of ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$ not only satisfies the Markov property but also remains invariant over time. That is, for any $t$, $a$ and $o$, we have $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a|\bm{H}_{t})=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\bm{O}_{t}})$ and $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\bm{O}_{t}}=o)=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\bm{A}_{1}}=a|{\color[rgb]{1,.5,0}\bm{O}_{1}}=o)$.
Following a given policy $\pi$, the expected cumulative reward the agent receives is given by $J(\pi)=\sum_{t\geq 0}\gamma^{t}\mathbb{E}^{\pi}({\color[rgb]{1,0,0}\bm{R}_{t}})$ where $10<\gamma<$ 1 denotes the discount factor that balances immediate and future rewards, and the expectation $\mathbb{E}^{\pi}$ is taken by selecting all actions according to $\pi$. The optimal policy $\pi^{*}$ is defined to be the maximizer of $J(\pi)$, i.e., $\pi^{*}=\arg\max_{\pi}J(\pi)$.
@@
    Markov assumption: At any time step $t$, given the current observation-action pair $({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})$, the current reward ${\color[rgb]{1,0,0}\bm{R}_{t}}$ and the next observation ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ are conditionally independent of the prior history $({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\bm{R}_{i}})_{i<t}$.
@@
    Stationarity assumption: For any observation $o$ and action $a$, the conditional distribution of ${\color[rgb]{1,0,0}\bm{R}_{t}}$ and ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ given ${\color[rgb]{1,.5,0}\bm{O}_{t}}=o$ and ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}=a$ does not depend on time $t$.
@@ 3 Hypothesis Testing in RL: Simulations and Case Studies
    High-order MDP (middle panel): In this model, the offline data does not satisfy the first-order Markov assumption but exhibits a high-order Markov property. Specifically, ${\color[rgb]{1,0,0}\bm{R}_{t}}$ and ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ depend not only on ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$ and ${\color[rgb]{1,.5,0}\bm{O}_{t}}$, but also on the past $k\geq 2$ lagged variables $({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\bm{R}_{i}})_{t-k<i<t}$. To apply existing RL algorithms, we merge the current observation ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ with these lagged variables to define a new state ${\color[rgb]{0,0,1}\bm{S}_{t}}$, so as to ensure the Markov property. The RL algorithm can then be applied to the transformed data triplets $({\color[rgb]{0,0,1}\bm{S}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}},{\color[rgb]{1,0,0}\bm{R}_{t}})_{t}$ to compute the optimal policy.
@@ 3.1 Case study: Diabetes datasets
We apply the hypothesis tests developed by Shi et al. (2020) and Zhou et al. (2023) [^4] Code available at: TestMDP and markov_test. to test whether the data follows a $k$-th order MDP for $k=1,2,\cdots,10$. Both tests assess the Markov assumption in RL. We will detail their methodologies in Section 4. The corresponding $p$-values are listed in Table 1. Small $p$-values indicate that the data does not conform to the $k$-th order MDP, whereas large $p$-values suggest that the null hypothesis that the data originates from a $k$-th order MDP cannot be rejected.
Table 1: $p$-values and expected returns of estimated optimal policies for the diabetes dataset, under MDP models across different orders, ranging from $k=1$ to $10$.
@@
Zhou et al. (2023)’s $p$-value 0 0.010 0.030 0.240 0.243 0.421 0.436 0.485 0.360 0.338
Shi et al. (2020)’s $p$-value 0 0.001 0.003 0.097 0.084 0.092 0.066 0.069 0.091 0.103
@@
At a significance level of $\alpha=0.05$, both tests by Shi et al. (2020) and Zhou et al. (2023) reject the null hypothesis for $k<4$ but do not reject it for $k\geq 4$. These results imply that the data likely comes from a 4-th order MDP, meaning that data collected from four hours earlier is conditionally independent of future observations, given the data from the current four hours. Additionally, it can be seen that the test by Shi et al. (2020) generally yields smaller $p$-values than those obtained from the test of Zhou et al. (2023), despite that both tests are consistent in theory. This discrepancy arises because they employ different methods to estimate the Markov transition function; see Section 4 for details. According to simulation results in Zhou et al. (2023), the test by Shi et al. (2020) sometimes suffers from inflated type-I errors. Consequently, $p$-values from Zhou et al. (2023) are more reliable.
To further demonstrate the utility of these tests in improving existing RL algorithms, we conduct the following cross-validation procedure: (i) Split the data into training and testing data subsets, each containing three patients data trajectories; (ii) Apply the fitted Q-iteration algorithm (FQI, Ernst et al., 2005; Riedmiller, 2005) to the training dataset to estimate the optimal policy under the $k$-th order MDP assumption; (iii) Apply the fitted Q-evaluation algorithm (FQE, Le et al., 2019) to the testing dataset to estimate the expected return of the resulting policy; (iv) Repeat this process across all train-test splits, average the results and report them in Table 1. The results clearly show that the policy estimated under the 4th-order MDP model assumption achieves the largest expected return, highlighting the importance of testing the Markov assumption in high-order MDPs. In particular, the Markov test can considerably enhance the performance of RL algorithms by enabling accurate determination of the MDP’s order.

### v2

@@ Abstract
Reinforcement learning (RL) is concerned with how intelligence agents take actions in a given environment to maximize the cumulative reward they receive. In healthcare, applying RL algorithms could assist patients in improving their health status. In ride-sharing platforms, applying RL algorithms could increase drivers’ income and customer satisfaction. For large language models, applying RL algorithms could align their outputs with human preferences. Over the past decade, RL has been arguably one of the most vibrant research frontiers in machine learning. Nevertheless, statistics as a field, as opposed to computer science, has only recently begun to engage with RL both in depth and in breadth. This chapter presents a selective review of statistical inferential tools for RL, covering both hypothesis testing and confidence interval construction. Our goal is to highlight the value of statistical inference in RL for both the statistics and machine learning communities, and to promote the broader application of classical statistical inference tools in this vibrant area of research.
@@ 1 Introduction
In the statistics literature, RL is closely related to a vast body of work on estimating dynamic treatment regimes (DTRs), which aims to identify personalized treatment strategies for individual patients to maximize their health outcomes; see Chakraborty and Moodie (2013); Kosorok and Moodie (2015); Kosorok and Laber (2019); Tsiatis et al. (2019); Li et al. (2023b) for reviews. However, most methods employ the so-called non-Markov decision process (Kallus and Uehara, 2020) – where rewards and transitions depend on the entire data history – to model the observed data. They often prove ineffective when applied to Markov decision processes (Kallus and Uehara, 2022) – which are more commonly studied in machine learning to formulate numerous sequential decision making problems. More recent work has shifted its focus to studying Markov decision processes for both policy optimization (Ertefaie and Strawderman, 2018; Luckett et al., 2020; Liao et al., 2022; Yang et al., 2022; Chen et al., 2024; Li et al., 2024; Shi et al., 2024a; Zhou et al., 2024) and policy evaluation (Liao et al., 2021; Hu and Wager, 2023; Ramprasad et al., 2023; Wang et al., 2023a; Shi et al., 2024b).
@@ 2 Background: Data, Policies, and Models in Reinforcement Learning
  • At each time step $t$, an agent observes certain features of the environment, denoted as ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$;
  • Based on ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$, the agent takes an action, denoted as ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$;
  • The environment provides a corresponding reward, denoted as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$, and transitions to a new state at the next time step.
This process repeats iteratively, as illustrated in Figure 1(a). Thus, the observed data can be summarized into a sequence of “observation-action-reward” triplets $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}})_{t\geq 0}$. It is worth noting that the observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ at each time step is not equivalent to the environment’s state ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{S}_{t}}$. Indeed, the state can be viewed as a special observation with the Markov property, and we will elaborate on the difference between the two later.
@@
    History-dependent policy: This is the most general form of policy. At each time $t$, we define $\bm{H}_{t}$ as the set containing the current observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ and all prior historical information $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{i}})_{i<t}$. Under this policy, the selection of ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$ depends on the entire history $\bm{H}_{t}$.
@@
    Markov policy: The second type is a subset of the first. Under this policy, the conditional distribution of ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$ depends only on the current observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$, rather than the entire history $\bm{H}_{t}$. In other words, for any $t$ and $a$, we have $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a|\bm{H}_{t})=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}})$.
@@
    Stationary policy: The third type is a subset of the second. Under this policy, the conditional distribution of ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$ not only satisfies the Markov property but also remains invariant over time. That is, for any $t$, $a$ and $o$, we have $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a|\bm{H}_{t})=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}})$ and $\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}=o)=\mathbb{P}({\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{1}}=a|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{1}}=o)$.
Following a given policy $\pi$, the expected cumulative reward the agent receives is given by $J(\pi)=\sum_{t\geq 0}\gamma^{t}\mathbb{E}^{\pi}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}})$ where $10<\gamma<$ 1 denotes the discount factor that balances immediate and future rewards, and the expectation $\mathbb{E}^{\pi}$ is taken by selecting all actions according to $\pi$. The optimal policy $\pi^{*}$ is defined to be the maximizer of $J(\pi)$, i.e., $\pi^{*}=\arg\max_{\pi}J(\pi)$.
@@
    Markov assumption: At any time step $t$, given the current observation-action pair $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})$, the current reward ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ and the next observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ are conditionally independent of the prior history $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{i}})_{i<t}$.
@@
    Stationarity assumption: For any observation $o$ and action $a$, the conditional distribution of ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ and ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ given ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}=o$ and ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}=a$ does not depend on time $t$.
@@ 3 Hypothesis Testing in RL: Simulations and Case Studies
    High-order MDP (middle panel): In this model, the offline data does not satisfy the first-order Markov assumption but exhibits a high-order Markov property. Specifically, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ and ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ depend not only on ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$ and ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$, but also on the past $k\geq 2$ lagged variables $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{i}})_{t-k<i<t}$. To apply existing RL algorithms, we merge the current observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ with these lagged variables to define a new state ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{S}_{t}}$, so as to ensure the Markov property. The RL algorithm can then be applied to the transformed data triplets $({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{S}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}})_{t}$ to compute the optimal policy.
@@ 3.1 Case study: Diabetes datasets
We apply the hypothesis tests developed by Shi et al. (2020b) and Zhou et al. (2023b) [^4] Code available at: TestMDP and markov_test. to test whether the data follows a $k$-th order MDP for $k=1,2,\cdots,10$. Both tests assess the Markov assumption in RL. We will detail their methodologies in Section 4. The corresponding $p$-values are listed in Table 3.1. Small $p$-values indicate that the data does not conform to the $k$-th order MDP, whereas large $p$-values suggest that the null hypothesis that the data originates from a $k$-th order MDP cannot be rejected.
Table 3.1: $p$-values and expected returns of estimated optimal policies for the diabetes dataset, under MDP models across different orders, ranging from $k=1$ to $10$.
@@
Zhou et al. (2023b)’s $p$-value 0 0.010 0.030 0.240 0.243 0.421 0.436 0.485 0.360 0.338
Shi et al. (2020b)’s $p$-value 0 0.001 0.003 0.097 0.084 0.092 0.066 0.069 0.091 0.103
@@
At a significance level of $\alpha=0.05$, both tests by Shi et al. (2020b) and Zhou et al. (2023b) reject the null hypothesis for $k<4$ but do not reject it for $k\geq 4$. These results imply that the data likely comes from a 4-th order MDP, meaning that data collected from four hours earlier is conditionally independent of future observations, given the data from the current four hours. Additionally, it can be seen that the test by Shi et al. (2020b) generally yields smaller $p$-values than those obtained from the test of Zhou et al. (2023b), despite that both tests are consistent in theory. This discrepancy arises because they employ different methods to estimate the Markov transition function; see Section 4 for details. According to simulation results in Zhou et al. (2023b), the test by Shi et al. (2020b) sometimes suffers from inflated type-I errors. Consequently, $p$-values from Zhou et al. (2023b) are more reliable.
To further demonstrate the utility of these tests in improving existing RL algorithms, we conduct the following cross-validation procedure: (i) Split the data into training and testing data subsets, each containing three patients’ data trajectories; (ii) Apply the fitted Q-iteration algorithm (FQI, Ernst et al., 2005; Riedmiller, 2005) to the training dataset to estimate the optimal policy under the $k$-th order MDP assumption; (iii) Apply the fitted Q-evaluation algorithm (FQE, Le et al., 2019) to the testing dataset to estimate the expected return of the resulting policy; (iv) Repeat this process across all train-test splits, average the results and report them in Table 3.1. The results clearly show that the policy estimated under the 4th-order MDP model assumption achieves the largest expected return, highlighting the importance of testing the Markov assumption in high-order MDPs. In particular, the Markov test can considerably enhance the performance of RL algorithms by enabling accurate determination of the MDP’s order.